OpenStudy (ja1):

What is the simplified form of :

OpenStudy (ja1):

\[\frac{ x^2 - 49 }{ x^2 - 14x + 49 }\]

OpenStudy (ja1):

Show steps please and/or explain how to do this, I am reviewing some material and I forgot how to do this.

OpenStudy (phi):

factor the top and bottom cancel any like terms

OpenStudy (phi):

the top is a "difference of squares" it factors this way a^2 - b^2 = (a-b)(a+b)

OpenStudy (ja1):

Ok so it would be: \[\frac{ (x - 7) (x + 7) }{ x^2 - 14x + 49 }\]

OpenStudy (phi):

for the bottom first list the pairs of factors of 49: 1,49 7,7 are all of them the + on +49 means the factors are the same sign the -14 from -14x means : the - means the largest factor is -. (so here both factors will be -) now look at the pairs. which pair, with both numbers being minus, add up to -14 those will be the factors

OpenStudy (ja1):

Well it would be -7 and -7 because if you add both those you get -14 but if you multiply two negatives you get a positive.

OpenStudy (ja1):

So now I would have: \[\frac{ (x + 7) (x-7) }{ (x-7) (x-7) }\]

OpenStudy (ja1):

Right?

OpenStudy (phi):

yes, now cancel (x-7) from top and bottom (because anything divided by itself is 1) also, to be complete, make a note: x≠7 (x is not allowed to by 7)

OpenStudy (ja1):

Oh ok so now I would get: \[\frac{ x + 7 }{ x - 7}\]

OpenStudy (ja1):

So that would be the simplified form since it cannot be simplified further right?

OpenStudy (phi):

yes, that is the answer

OpenStudy (ja1):

Woohoo :D Thanks a bunch, I had forgotten how to do this